HIGH-RESOLUTION MEASUREMENTS OF THE BROADENING AND SHIFTING OF CO-N_{2} AT LOW-TEMPERATURES

\underline{RICHARD WEHR}, A. VITCU, R. CIURYLO, J. R. DRUMMOND and A. D. MAY, Department of Physics, University of Toronto, ON M5S 1A7, Canada.

Atmospheric remote sounding requires an increasingly accurate knowledge of the spectral properties of gases under environmental conditions. To that end, high-resolution measurements of the broadening and shifting of several CO-N_{2} fundamental P and R branch lines are presented at −45^{∘}C and −10^{∘}C, and pressures of about 50 kPa. Combining our broadening results with previously published data ^{a}, the measured temperature dependence of the broadening coefficients is compared with the temperature dependence predicted by the usual scaling law, γ_{1}/γ_{2} = (T_{1}/T_{2})^{−n}. It is shown that the law is no longer accurate within experimental error. The temperature dependence of the shifts is also measured.

aA. Predoi-Cross, C. Luo, P. M. Sinclair, J. R. Drummond, and A. D. May, J. Mol. Spec. 198, 291-303 (1999).